OpenStudy (anonymous):

The area of a rectangle is 20x^10. The length of one side is 4x^2. What is the width?

OpenStudy (anonymous):

5x^8, we know a=b*h thus 20x^10=(4x^2)b(x), where b(x) is a function for the base, simplifying g(x)=5x^8